# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v2

@@ 4 Simulation Study
In this section, we carry out extensive simulations of both divide-and-conquer algorithms. The goals of this section are to compare the performance of the algorithms under different sub-sampling schemes and see if there exists an optimal sub-sampling method. Additionally, we also hope to empirically validate the CP theoretical results. All experiments in this and the subsequent section were carried out in R on a MacBook Pro with Apple M3 chip, 16 GB of memory parallelized using 10 cores. All code is available on the author’s GitHub: https://github.com/eyanchenko/subsamp.
@@ 4.1.1 Settings
There are six simulations settings. In each, we generate networks, apply Algorithm 1 and compare the estimated membership vector with the true labels using the adjusted rand index (ARI). ARI provides a measure of concordance between true and estimated labels with 1 being the largest value and indicating a perfect match (up to permuting the label names). We repeat this process for 100 MC iterations and record the average ARI. [^3] The run-time was also recorded but we do not report it. As some sub-sampling routines used optimized functions from R packages while others are written by hand, it is not a fair comparison. In general, the exploration-based sub-sampling methods will be slower.
We generate networks with a community structure by using a $K=2$-block SBM. Recall from Section 3, if $A$ is the corresponding adjacency matrix with data-generating model $P$, then $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(P_{ij})$ where $P_{ij}=p_{c_{i},c_{j}}$. We take $p_{11}=p_{22}>p_{12}$ to generate networks with assortative community structure where $p_{11}=p_{22}$ and $p_{12}$ represent the intra-community and inter-community edge probabilities, respectively. We set $p_{12}=0.01$ for all experiments in this sub-section.
An overview of the simulation settings is available in Table 2. Setting 1 looks at the effect of the intra-community edge probability where larger values indicated stronger community structure. Settings 2 and 3 increase the number of nodes in the network with $B$ held fixed and increasing with $n$, respectively. The proportion of nodes in one community is the subject of setting 4 where greater imbalance in the community sizes should make the task more difficult. Finally, settings 5 and 6 look at the hyper-parameters in the $\mathsf{PACE}$ model, namely the number of sub-graphs drawn and size of sub-graphs, respectively. For all experiments, we use the fast greedy algorithm (Clauset et al.,, 2004) as our base community detection algorithm as in Zhang et al., (2022).
Setting $n$ $p_{11}$ $\kappa_{1}$ $B$ $qn$
@@
5 $5000$ $0.04$ $0.75$ ${\bf(100,10,000)}$ $250$
6 $5000$ $0.04$ $0.75$ $1000$ ${\bf(100,500)}$
Table 2: Simulation settings for $\mathsf{PACE}$ divide-and-conquer algorithm. $n$ is the number of nodes in the network; $p_{11}$ is the intra-community edge probability; $\kappa_{1}$ is the proportion of nodes in community 1; $B$ is the number of sub-graphs sampled; and $qn$ is the number of nodes per sub-graph. The parameters which vary are highlighted in ${\bf bold}$.
@@ 4.1.2 Results
The results are in Figure 1. In Setting 1, save BFS and DFS, all methods yield an increasing ARI as $p_{11}$ increases. When $p_{11}\leq 0.04$, RNN has the largest ARI while RN does the best when $p_{11}>0.04$. BFS improves at first before quickly dropping off, while DFS yields extremely low ARI values. Indeed, in all settings, it is clear that DFS performs poorly. Turning to setting 2, most methods’ ARI increases with $n$ until $n=5000$ before decreasing. By $n=9000$, these methods experience a steep drop off in performance. When $B$ also increases with $n$ as in setting 3, DN, RE, RN, RNN and RW all have increasing ARI, with no clear distinction.
The effects of the distribution of nodes between communities is evident in setting 4. Indeed, as the nodes are more disproportionately allocated between groups, each method decreases in performance with RN being the most negatively affected. As $B$ increases in setting 5, all methods, save DFS, have an improvement in performance. Interestingly, while RN and RNN plateau at an ARI of 1, DN, RE, RW and BFS all begin to yield slightly decreasing values of ARI for $B>2500$. Finally, setting 6 shows that as the number of nodes in the sub-samples increases, the communities are easier to detect for all methods except BFS and DFS. RN has the largest ARI followed by RNN for most values of $q$.
@@ 4.1.3 Discussion
From the results of these simulations, it appears that RN and RNN perform the best with RW, RE and DN also yielding good results. Even more clearly, the results demonstrate that BFS and DFS are poor samplers to use with the $\mathsf{PACE}$ algorithm. Setting 2 reveals the importance of increasing the number of sub-samples as $n$ increases, even though the theoretical results depend on a fixed $B$. Settings 5 and 6 show that $B$ and $q$ must be carefully chosen, where small values can severely hinder performance. Finally, all methods suffer when the nodes are unequally distributed between communities. This is not surprising as the identification problem is much more challenging in this setting. Indeed, the presence of small communities likely requires the increase in the number of nodes per sub-sample in order to capture the structure.
Figure 1: Community detection simulation results. The number in the upper-left corner corresponds to the simulation setting (1-6).
@@ 4.2.1 Settings
We again look at six simulation settings. For each setting, we generate a network, apply Algorithm 2 and compute the AUC between the outputted proportions and true labels. We vary different aspects of the data-generating model as well as the algorithm hyper-parameters. All networks are generated from a CP-SBM. The networks are generated the same as in the community detection simulations except now $p_{11}>p_{12}>p_{22}$ where $p_{11},p_{12},p_{22}$ represent the core-core, core-periphery, and periphery-periphery edge probabilities, respectively. Each setting is repeated for 100 MC replications and the average AUC is reported.
Setting $n$ $p_{11}$ $\alpha_{n}$ $B$ $qn$
@@
Table 3: Simulation settings for CP divide-and-conquer algorithm. $n$ is the number of nodes in the network; $p_{11}$ is the core-core edge probability; $\alpha_{n}$ is the proportion of nodes in the core; $B$ is the number of sub-graphs sampled; and $qn$ is the number of nodes per sub-graph. The parameters which vary are highlighted in ${\bf bold}$.
The simulation settings are summarized in Table 3. For each setting, we fix $p_{22}=0.001$ and set $p_{12}=\frac{1}{2}p_{11}$. In setting 7, we vary $p_{11}$ where larger values correspond to a greater strength of a CP structure. Settings 8 and 9 look at the effect of increasing the network size where the former fixes $B$ while the latter allows $B$ to increase with $n$. The effect of the core size is studied in setting 10. Choice of hyper-parameters are investigated in settings 11 and 12, varying the values of $B$ and $q$, respectively.
@@ 4.2.2 Results
The results are in Figure 2. In setting 7, as $p_{11}$ increases, all sub-sampling schemes have an increasing AUC. BFS, DN, RE, and RW all perform nearly identically with DFS also performing the same once $p_{11}\geq 0.005$. On the other hand, RN and RNN clearly yield worse results. Next, it is helpful to compare settings 8 and 9 together since they both look at increasing $n$. Similar to setting 7, DFS yields the best results with BFS, DN, RE, and RW yielding similar performance, while RN and RNN consistently yield the lowest AUC values. In addition, we notice that if $B$ is fixed and $n$ increases (setting 8), the AUC values start to decrease for larger $n$, most notably for RN and RNN. In setting 9, however, where $B$ increases with $n$, all methods save RNN have generally monotonically increasing AUC with increasing $n$.
Moving to setting 10, the performance of DFS drastically decreases as the core size increases. All other methods have roughly constant AUC values. In setting 11, we see that increasing the number of sub-sampled nodes beyond one or two thousand yields minimal returns for DFS, BFS, DN, RE, and RW. RN, however, has monotonically improving AUC with increasing $B$ while RNN also improves with larger $B$. These trends are similar in setting 12. Increasing the number of nodes in the sub-samples only seems to noticeably improve RN and RNN; the other methods remain relatively unchanged, especially once $qn>50$.
Figure 2: Core-periphery simulation results. The number in the upper-left corner corresponds to the simulation setting (7-12).
@@ 4.2.3 Discussion
There are several takeaways from these simulation settings. First, it is clear that BFS, DN, RE, and RW are the best sub-sampling methods for CP identification. DFS also performs well but its behavior for increasing core size is somewhat concerning. On the other hand, RN and RNN display the worst performance. Since the mis-classification rate is lower when core nodes are sampled more frequently, the best methods will have a higher probability of sampling core nodes. Indeed, both RN and RNN sample core nodes less frequently than BFS, DN, RE, and RW. This also aligns with the theoretical results which showed that RE and DN should perform similarly, and that both of these methods should outperform RN. Continuing on a theoretical thread, we note that the theory in Section 3 assumes that the number of sub-graphs, $B$, is fixed. In Setting 8, however, we saw that this assumption can lead to poor performance empirically. Indeed, settings 8 and 9 seem to indicate that $B$ must be increased in larger networks, as there are more nodes to sample. Lastly, these results have implications on the choice of hyper-parameters. In settings 11 and 12, we found that increasing $B$ and $q$ only leads to marginal gains in performance, particularly for BFS, DN, RE, and RW. This means that these values can be set relatively small, increasing the computational speed. RN and RNN require larger $B$ and $q$ to match the other method’s performance, further demonstrating their weakness for this task.
@@ 5 Real-data analysis
In this section, we implement both divide-and-conquer algorithms on various real-world networks.
@@ 5.1 Community detection
We apply the $\mathsf{PACE}$ algorithm to two different real-world networks. This first is Political Blogs (Adamic and Glance,, 2005) ($n=1490,m=16,715$) where nodes represent blogs and edges arise if one blog references another. The blogs are divided between liberal and conservative, so we set $K=2$, and keep only the largest connected component, yielding $n=1222$. The second is Facebook (Leskovec and Mcauley,, 2012) ($n=4039,m=88,234$) where nodes are users connected with an edge if they are friends. This is an ego network with ten egos so we set $K=10$. We remove all edge weights and directions as well as self-loops for both networks.
We apply the $\mathsf{PACE}$ algorithm to each network and report the size of the largest community as well as the modularity of the returned node membership vector where a large value indicates a better community assignment. Since we have the ground truth labels (which we call the Oracle) for the Political Blogs network, we can also compute the ARI between the true and estimated labels. Finally, we set $q=250/n$ for each network, and $B=1000$ and $5000$ for Political Blogs and Facebook, respectively.
The results are in Table 4. For Political Blogs, RN yields both the largest modularity value as well as most similar labels to the ground-truth. RNN does the next best, with RE also performing similarly. On the other hand, DFS yields substantially lower modularity and ARI values, likely due to its much greater imbalance in community labels. The results are similar for Facebook, where RN has the largest modularity and RNN provides the second-largest. In this example, both BFS and DFS provided imbalanced class labels, leading to lower modularity values.
@@
Method Size $Q$ ARI Size $Q$
Oracle 0.52 0.405
RN 0.53 0.425 0.810 0.46 0.743
DN 0.63 0.403 0.430 0.57 0.625
RE 0.66 0.414 0.495 0.64 0.550
BFS 0.62 0.388 0.463 0.86 0.041
DFS 0.90 0.075 0.061 0.99 0.003
RNN 0.63 0.419 0.584 0.66 0.694
RW 0.62 0.406 0.454 0.72 0.546
Table 4: Real data analysis results for PACE community structure divide and conquer algorithm.

### v3

@@ 4 Simulation Study
In this section, we carry out extensive simulations of both divide-and-conquer algorithms. The goals of this section are to compare the performance of the algorithms under different sub-sampling schemes, as well as compare the divide-and-conquer algorithms against the original algorithm applied to the entire graph. Additionally, we hope to empirically validate the CP theoretical results. All experiments in this section were carried out in R on a Mac Mini with Apple M4 chip and 16 GB of memory parallelized using 9 cores. The code to replicate these simulations is available on the author’s GitHub: https://github.com/eyanchenko/subsamp.
@@ 4.1.1 Settings
There are six simulations settings. In each, we generate networks, apply Algorithm 1 and compare the estimated membership vector with the true labels using the adjusted Rand index (ARI) (Rand,, 1971; Hubert and Arabie,, 1985). The Rand index, $R(\hat{\boldsymbol{c}},\boldsymbol{c})$, provides a measure of concordance between the true, $\boldsymbol{c}$, and estimated communities, $\hat{\boldsymbol{c}}$. We define $c_{i}=k$ if node $i$ is in community $k$ for $k\in\{1,\dots,K\}$, and similarly, $\hat{c}_{i}=k$ if Algorithm 1 assigns node $i$ to community $k$. The Rand index is then computed as
$R(\hat{\boldsymbol{c}},\hat{\boldsymbol{c}})=\frac{1}{{n\choose 2}}\sum_{i<j}^{n}\gamma_{ij}(\hat{\boldsymbol{c}},\boldsymbol{c})$
where
$\gamma_{ij}(\hat{\boldsymbol{c}},\boldsymbol{c})=\begin{cases}1,&\hat{c}_{i}=\hat{c}_{j}\text{ and }c_{i}=c_{j}\\ 1,&\hat{c}_{i}\neq\hat{c}_{j}\text{ and }c_{i}\neq c_{j}\\ 0,&\text{otherwise}.\end{cases}$
The adjustment in the ARI accounts for the fact that even a poor community detection algorithm could correctly classify some nodes just by chance. The maximum value of the ARI is 1 which indicates a perfect match between the estimated and true community labels (up to permutation of community labels).
We generate networks with a community structure by using a $K=2$-block SBM. Recall from Section 3, if ${\bf A}$ is the corresponding adjacency matrix with data-generating model ${\bf P}$, then $A_{ij}\stackrel{{\scriptstyle ind.}}{{\sim}}\mathsf{Bernoulli}(P_{ij})$ where $P_{ij}=p_{c_{i},c_{j}}$. We take $p_{11}=p_{22}>p_{12}$ to generate networks with assortative community structure where $p_{11}=p_{22}$ and $p_{12}$ represent the intra-community and inter-community edge probabilities, respectively. We set $p_{12}=0.01$ for all experiments in this sub-section.
[…]
Setting $n$ $p_{11}$ $\alpha$ $B$ $qn$
@@
[…]
6 $5000$ $0.05$ $0.75$ $1000$ ${\bf(100,500)}$
[…]
@@ 4.1.2 Results
The results are in Figures 1 and 2. In setting 1, save BFS and DFS, all methods yield an increasing ARI as $p_{11}$ increases. When $p_{11}\leq 0.04$, RNN has the largest ARI while RN does the best when $p_{11}>0.04$ among the divide-and-conquer methods. BFS improves at first before quickly dropping off, while DFS yields extremely low ARI values. Indeed, in all settings, it is clear that DFS performs poorly. Moreover, the algorithm applied to the full network yields the best performance when $p_{11}>0.03$. Turning to setting 2, the ARI increases with $n$ until around $n=4000$ before decreasing for most methods. By $n=8000$, these methods have experienced a drop off in performance. The full algorithm, on the other hand, has monotonically increasing performance with $n$, and yields the largest ARI values when $n\geq 4000$. When $B$ also increases with $n$ as in setting 3, DN, RE, RN, RNN and RW all have increasing ARI. RN and RNN perform slightly better and exceed or essentially match the ARI of Full for all $n$
The effects of the distribution of nodes between communities is evident in setting 4. Indeed, as the nodes are more disproportionately allocated between groups, each method decreases in performance with RN and Full being the most severely affected. As $B$ increases in setting 5, all methods, save DFS, see an improvement in performance. Interestingly, while RN and RNN plateau at an ARI of 1, DN, RE, RW and BFS all begin to yield slightly decreasing values of ARI for $B>2500$. Finally, setting 6 shows that as the number of nodes in the sub-samples increases, the communities are easier to detect for all methods except BFS and DFS. Among the divide-and-conquer methods, for most values of $q$, RN has the largest ARI followed by RNN. Note that the results for Full are constant in the last two settings since $B$ and $q$ do not appear as parameters in the full algorithm.
The computation time for each approach is in Figure 2. Somewhat surprisingly, Full yields the fastest computation time across all settings. The different $\mathsf{PACE}$ algorithms have similar computation times, with the exception of RNN, which consistently yields the slowest times. This is likely because this sub-sampling algorithm was written by hand, while the other sub-samplers were implemented from the igraph package (Csárdi et al.,, 2025).
@@ 4.1.3 Discussion
The first major takeaway from the simulation study is the strong performance of the base algorithm applied to the entire network. Not only did it yield the best detection results across many settings (as we might have expected), but it was also the fastest algorithm as well. The advantage of the divide-and-conquer approach should be that it improves computation time. If the base algorithm is fast enough, as in this case, then evidently it does not. In Mukherjee et al., (2021), the authors primarily discuss using spectral clustering or likelihood-based method as the base algorithm, which seem to be slower than Fast Greedy. In this case, $\mathsf{PACE}$ will yield computational advantages. For example, in the Supplemental Materials where we use Walktrap as the base algorithm, the divide-and-conquer approaches are indeed faster. This implies that depending on the choice of base algorithm, the advantages of $\mathsf{PACE}$ may differ. Even though $\mathsf{PACE}$ is slower, there are several settings where it yields larger ARI values, in particular for small $p_{11}$ (setting 1), small $n$ (settings 2 and 3) and highly disproportionate communities (setting 4). The similarity between these settings is that the community structure is fairly weak, either because $p_{11}$ and $p_{12}$ are close, the network is small, or one community dominates the network. In such cases, it could be that the sub-sampling approach is averaging out some of the noise in the network, yielding better identification results.
Looking only at the divide-and-conquer approaches, it appears that RN and RNN perform the best with RW, RE and DN also yielding good results. Even more clearly, the results demonstrate that BFS and DFS are poor samplers to use with the $\mathsf{PACE}$ algorithm. This is likely because the sub-graphs resulting from them are dissimilar from the original graph and do not preserve the community structure. This result is in agreement with Krishnamurthy et al., (2005) which showed poor performance by both exploration samplers. Setting 2 reveals the importance of increasing the number of sub-samples as $n$ increases, even though the theoretical results depend on a fixed $B$. Settings 5 and 6 show that $B$ and $q$ must be carefully chosen since small values can severely hinder performance. Finally, all methods suffer when the nodes are unequally distributed between communities. This is not surprising as the identification problem is much more challenging in this setting. Indeed, the presence of small communities likely requires increasing in the number of nodes per sub-sample in order to capture this structure. In the Supplemental Materials, we show that conclusions are similar when using Walktrap as the base community detection algorithm.
Figure 1: Community detection simulation results with Fast Greedy base algorithm. The number in the upper-left corner corresponds to the simulation setting (1-6).
Figure 2: Community detection simulation run-time results with Fast Greedy base algorithm. The number in the upper-left corner corresponds to the simulation setting (1-6).
@@ 4.2.1 Settings
We again look at six simulation settings. For each setting, we generate a network, apply Algorithm 2 and compute the area under the receiver operator curve (AUC) between the outputted proportions and true labels. CP identification is fundamentally about classifying nodes as core or periphery and AUC is a popular metric for such tasks. The base CP detection algorithm is a greedy algorithm based on Algorithm 1 from Yanchenko, (2022). Similar to the community detection simulations, we also apply this algorithm to the entire network as a benchmark for comparison (Full). The networks are generated the same way as in the community detection simulations except now $p_{11}>p_{12}>p_{22}$ where $p_{11},p_{12},p_{22}$ represent the core-core, core-periphery, and periphery-periphery edge probabilities, respectively. For each setting, we fix $p_{22}=0.001$ and set $p_{12}=\frac{1}{2}p_{11}$.
Setting $n$ $p_{11}$ $\alpha$ $B$ $qn$
@@
[…]
@@ 4.2.2 Results
The results are in Figure 3. In setting 7, as $p_{11}$ increases, all sub-sampling schemes have a non-decreasing AUC. BFS, DN, RE and RW all perform nearly identically with DFS performing the same once $p_{11}\geq 0.005$. On the other hand, RN and RNN yield worse performance, while Full’s AUC plateaus before reaching 1. Next, it is helpful to compare settings 8 and 9 together since they both look at increasing $n$. Similar to setting 7, DFS yields the best results with BFS, DN, RE and RW yielding similar performance. RN and RNN consistently yield the lowest AUC values in both settings. In addition, we notice that if $B$ is fixed and $n$ increases (setting 8), the AUC values start to decrease for larger $n$, most notably for RN and RNN. In setting 9, however, where $B$ increases with $n$, all methods save RNN have a generally monotonically increasing AUC with increasing $n$. In both of these settings, the divide-and-conquer algorithms tend to outperform Full.
Moving to setting 10, the performance of DFS noticeably decreases as the core size increases. All other methods have roughly constant AUC values. In setting 11, we see that increasing the number of sub-sampled nodes beyond one or two thousand yields minimal returns for DFS, BFS, DN, RE and RW. RN, however, has monotonically improving AUC with increasing $B$ while RNN also improves with larger $B$. These trends are similar in setting 12. Increasing the number of nodes in the sub-samples only seems to noticeably improve RN and RNN; the other methods remain relatively unchanged, especially once $qn>50$.
We also report the computation time for each approach in Figure 4. In settings 7, 8, 9, and 10, all methods are significantly faster than the full algorithm (save RNN in 9). From setting 11, if the number of sub-samples is large, then the divide-and-conquer algorithm’s computation time is comparable to that of the full algorithm. RNN again consistently yields the slowest times for the divide-and-conquer approaches.
Figure 3: Core-periphery simulation accuracy results. The number in the upper-left corner corresponds to the simulation setting (7-12).
Figure 4: Core-periphery simulation run-time results. The number in the upper-left corner corresponds to the simulation setting (7-12).
@@ 4.2.3 Discussion
There are several takeaways from these simulations. Surprisingly, the divide-and-conquer algorithms outperform the original algorithm applied to the full network, both in terms of AUC and computing time. We expected there to be computational benefits to the sub-sampling scheme, but the superior AUC is somewhat unexpected. We postulate that this is because the sub-sampling step is able to de-noise the results. By averaging the CP results over many sub-graphs, it seems that any outlying behavior of nodes on the original graph can be removed and/or a weak CP structure can be detected. Next, it is clear that BFS, DN, RE and RW are the best sub-sampling methods for CP identification. DFS also performs well but its behavior for increasing core size is somewhat concerning. On the other hand, RN and RNN display the worst performance. Since the mis-classification rate is lower when core nodes are sampled more frequently, the best methods will have a higher probability of sampling core nodes. Indeed, both RN and RNN sample core nodes less frequently than BFS, DN, RE and RW. This also aligns with the theoretical results which showed that RE and DN should perform similarly, and that both of these methods should outperform RN. Additionally, RW and RE results are almost indistinguishable across all settings, as expected based on Lovász, (1993). Continuing on the theoretical thread, we note that the theory in Section 3 assumes that the number of sub-graphs, $B$, is fixed. In setting 8, however, we saw that this assumption can lead to poor performance empirically. Indeed, settings 8 and 9 seem to indicate that $B$ must be increased in larger networks, as there are more nodes to sample. Lastly, these results have implications on the choice of hyper-parameters. In settings 11 and 12, we found that increasing $B$ and $q$ only leads to marginal gains in performance, particularly for BFS, DN, RE and RW. This means that these values can be set relatively low, increasing the computational speed. RN and RNN require larger $B$ and $q$ to match the other method’s performance, further demonstrating their weakness for this task.
@@ 5 Real-data analysis
In this section, we consider real-world networks from various domains to study the performance of the algorithms under different settings and types of network. The goal is to understand how the choice of sub-sampler affects the results of the divide-and-conquer algorithms. Moreover, we are interested in considering the theoretical results on non-synthetic data.
@@ 5.1 Community detection
[…]
We apply the $\mathsf{PACE}$ algorithm to each network using Fast Greedy (Clauset et al.,, 2004) as the base community detection algorithm. We report the proportion of nodes in the largest community, $\hat{\alpha}$, as well as the modularity of the returned node membership vector, $Q$ (Newman,, 2006). The modularity is a measure of the strength of community structure in a network where a large value indicates a better community assignment. For a network ${\bf A}$ and community labels $\boldsymbol{c}$, it is computed as
$Q({\bf A},\boldsymbol{c})=\frac{1}{m}\sum_{i<j}\left(A_{ij}-\frac{d_{i}d_{j}}{2m}\right)\mathbb{I}(c_{i}=c_{j})$
where $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$. Since we have the ground truth labels (which we call the Oracle) for the Political Blogs network, we can also compute the ARI between the true and estimated labels. Finally, we set $q=250/n$ for each network, and $B=1000$ and $5000$ for Political Blogs and Facebook, respectively.
The results are in Table 3. For Political Blogs, RN yields both the largest modularity as well as most similar labels to the ground-truth. RNN does the next best, with RE also performing similarly. On the other hand, DFS yields substantially lower modularity and ARI values, likely due to its much greater imbalance in community labels (large $\hat{\alpha}$). The results are similar for Facebook, where RN has the largest modularity and RNN provides the second-largest. In this example, both BFS and DFS provided imbalanced class labels, leading to lower modularity values.
@@
Method $\hat{\alpha}$ $Q$ ARI $\hat{\alpha}$ $Q$
Oracle 0.52 0.41
Full 0.52 0.43 0.78 0.24 0.78
RN 0.53 0.42 0.81 0.46 0.74
DN 0.63 0.40 0.43 0.57 0.63
RE 0.66 0.41 0.50 0.64 0.55
BFS 0.62 0.39 0.46 0.86 0.04
DFS 0.90 0.08 0.06 0.99 0.00
RNN 0.63 0.42 0.58 0.66 0.69
RW 0.62 0.41 0.45 0.72 0.55
Table 3: Real data analysis results for PACE community structure divide and conquer algorithm. $\hat{\alpha}$ is the proportion of nodes in the largest community, $Q$ is the modularity and ARI is the adjusted Rand index between the estimated and ground-truth community labels.